OpenStudy (anonymous):

Given that cosΘ = 1/2 and that Θ lies in quadrant IV, determine the value of tanΘ.

OpenStudy (anonymous):

find \(\sin(\theta)\) then tangent is sine over cosine

OpenStudy (anonymous):

how do I find sinΘ?

OpenStudy (anonymous):

you can find that point on the last page of this cheat sheet look in quadrant IV where the first coordinate is \(\frac{1}{2}\) and the sine will be the second coordinate there are other ways to do it, but it is easy to see on the sheet

OpenStudy (anonymous):

the other way is to compute \[-\sqrt{1-\left(\frac{1}{2}\right)^2}\]

OpenStudy (dayakar):



OpenStudy (anonymous):

-√(1+(1/2)^2) = -√5/2

OpenStudy (anonymous):

minus, not plus

OpenStudy (anonymous):

\[\huge -\sqrt{1\color{red}-\left(\frac{1}{2}\right)^2}\]

OpenStudy (dayakar):

AB^2 = OB^2 - OA^2 = 2^2 - 1^2 = 4-1 =3 \[AB=-\sqrt{3}\]

OpenStudy (anonymous):

unit circle cheat sheet is very helpful here

OpenStudy (dayakar):

\[TAN \theta = \frac{ AB }{ OA }\]

OpenStudy (anonymous):

Thanks!